OpenStudy (anonymous):

Solve (x + 2)^2 + 7(x + 2) = 18. Identify any double roots. A. {–11, 0} B. {–2, –7} C. {–9, 2 (double root)} D. {–2, 9 (double root)}

OpenStudy (anonymous):

Do I simplify first or?...

OpenStudy (anonymous):

Yeah simplify by finding that \[(x+2)^{2} = x ^{2}+4x+4\] Multiply out the 7(x+2) part to get 7x+14 add like terms so you get \[x ^{2}+4x+4+7x+4 = 18\] You get\[x ^{2}+11x+18=18\] Subtract 18 from both sides:\[x ^{2}+11x = 0\]Factor an x out you find you get \[x(x+11) = 0\] set both x = 0 and (x+11) = 0 to find x = 0 and x = -11 Answer is: A.

OpenStudy (anonymous):

Thank you so much!!!

OpenStudy (anonymous):

No problem.